Clara bought 1 sandwich and 2 sodas for $6.25. A sandwhich cost $1 more than a soda. How much does one sandwhich cost?
Answer by Aldorozos(172)

Let's say price of a sandwich is S and price of a soda is a
S+2a = 6.25
S=1+a
We replace S in the first equation with 1+a
1+a +2a = 6.25 therefore 3a = 5.25 this means a = soda = 1.75
Sandwish is a dollar more which is 2.75